OpenStudy (anonymous):

i need help with factoring i apologize but i just dont understand factoring at all, i have a word problem question and it would really be appreciated if someone could help "Using complete sentences, explain how to factor each one. Be sure that the final factorization (or "answer") is a part of your explanation. 2x2 + 13x + 15 2x2 + 27x + 13" any help would be appreciated.

jhonyy9 (jhonyy9):

opinion ?

jhonyy9 (jhonyy9):

2x^2 +10x +3x +15 = (2x^2 +10x) +(3x +15) =2x(x+5) +3(x+5) = (2x +3)(x+5) so this is understandably sure now ?

jhonyy9 (jhonyy9):

so than make the second the same

zepdrix (zepdrix):

 This is another method you could try shadow :D It's more based on trial and error, and noticing a few little things about the equation. Probably not as good as factoring by grouping, but maybe it will help you understand. :)